OpenStudy (sphott51):

Solving equations - What is the value of x if (h+5)/x - 3 = 12? All I know the answer is x= (h) / 15 + 1/3.. I just need to do the work.

OpenStudy (anonymous):

any idea how to start this?

OpenStudy (sphott51):

No

OpenStudy (anonymous):

ok solets get rid of the fraction (h+5)/x - 3 = 12 h+5 = 12(x - 3) understand what i did? next expand the brackets for me :)

OpenStudy (anonymous):

\[\frac{ h+5 }{ x - 3 } = 12\]

OpenStudy (anonymous):

So that's the equation o: Now..let's put in what Hawk said: \[h + 5 = 12(x -3)\]

OpenStudy (anonymous):

Use the distributive property; okay?

OpenStudy (sphott51):